OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (2.4, 3) and (5, y)

jimthompson5910 (jim_thompson5910):

because we have an inverse variation, this means that each pair of points multiply to the same number k so, 2.4*3 = k 5*y =k and we can also say that 2.4*3 = 5y after doing a substitution

jimthompson5910 (jim_thompson5910):

so from here you solve 2.4*3 = 5y for y tell me what you get

OpenStudy (anonymous):

y = 1.44 thanks for the help

jimthompson5910 (jim_thompson5910):

correct